OpenStudy (anonymous):

Factor completely: 3x^2 + 5x + 1

OpenStudy (anonymous):

Can it be factored?

OpenStudy (anonymous):

I don't think so because there aren't any like terms?

OpenStudy (anonymous):

This one doesn't factor, as far as I can tell. Using the ac method, there are no factorizations of three that add to five.

mathslover (mathslover):

\[\huge{x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}}\] \[\huge{x = \frac{-5\pm\sqrt{25-12}}{2(3)}}\] \[\huge{x=\frac{-5\pm\sqrt{13}}{6}}\] the solutions are : \[\Huge{x_1=\frac{-5+\sqrt{13}}{6}}\] \[\Huge{x_2=\frac{-5-\sqrt{13}}{6}}\]

mathslover (mathslover):

wait

OpenStudy (anonymous):

Not exactly a factorization yet...

mathslover (mathslover):

wait

OpenStudy (anonymous):

Wait, so it would be prime right?

mathslover (mathslover):

this means that : \[\Huge{x-(\frac{-5+\sqrt{13}}{6})}\] \[\Huge{x-(\frac{-5-\sqrt{13}}{6})}\]

mathslover (mathslover):

these are the factors

OpenStudy (anonymous):

Do y'all need the answer choices?

OpenStudy (anonymous):

That would help.

OpenStudy (anonymous):

It could be shown as\[(x-x_1)(x-x_2)=3x^2+5x+1\]With x_1 and x_2 as above.

OpenStudy (anonymous):

It might be considered prime with respect to the integers or rationals, since the roots are irrational.

OpenStudy (anonymous):

A. (3x + 1)(x + 1) B. (3x + 5)(x + 1) C. (3x − 5)(x + 1) D. Prime

mathslover (mathslover):

Prime

OpenStudy (anonymous):

D is the only answer of those that makes sense.

OpenStudy (anonymous):

Yup D :).

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome!